OpenStudy (anonymous):

if a^2+b^2+c^2 = 2(a-b-c)-3, then (a-b+c)=? a) -1 b) 3 c) 1 d) -2

OpenStudy (anonymous):

actually given expression is a sum of three squares

OpenStudy (anonymous):

ok\[a^2+b^2+c^2 = 2(a-b-c)-3\]\[a^2+b^2+c^2-2a+2b+2c+1+1+1=\]\[\]

OpenStudy (anonymous):

(a-1)^2+(b+1)^2+(c+1)^2=0

OpenStudy (anonymous):

so a=1, b=-1, c=-1

OpenStudy (anonymous):

so 1-b+c=1

OpenStudy (anonymous):

a-b+c=1